# Wavehax: Aliasing-Free Neural Waveform Synthesis Based on 2D Convolution and Harmonic Prior for Reliable Complex Spectrogram Estimation

### v1

@@ II-D Time-frequency domain neural vocoders
HiNet [46] hierarchically estimates the log amplitude and phase spectrogram, ultimately reconstructing the waveform using inverse STFT (iSTFT). It estimates log amplitude spectrogram using a feed-forward network and conditions a neural source-filter (NSF) [47] model with the log amplitude spectrogram to generate a time-domain waveform, from which the phase spectrogram is extracted via STFT. The excitation generation module of the NSF constructs harmonic signals by accumulating phase rotation angles for each time step, determined by $\text{F}_{0}$. This process ensures a coherent phase structure of the output waveforms, aiding efficient phase spectrogram estimation and allowing HiNet to sidestep the challenges of direct phase estimation. However, the phase spectrogram derived from time-domain processing involving nonlinear operations can introduce aliasing.
@@ III-A1 Time-domain convolution
which can be implemented as a 1D convolutional layer in neural networks. To simplify the equations, we assume periodic padding, although zero padding or reflection padding can also be applied. According to the convolution theorem, this time-domain convolution modifies each frequency component $\mathcal{F}\{\bm{x}\}[k]$ depending on $\mathcal{F}\{\bm{h}\}$ in the frequency domain. Thus, it operates as a linear system with a frequency response characterized by the magnitude |$|\mathcal{F}\{\tilde{\bm{h}}\}|$ and phase $\angle\mathcal{F}\{\tilde{\bm{h}}\}$.
@@ III-B2 Equivalent interpretation and implementation
Building on the discussion in Section III-A2, we can derive an alternative implementation for computing $f(\hat{\bm{x}})$ in Eq. (7) using a coefficient signal $\hat{\bm{a}}\in\mathbb{R}^{2N}$ as follows:
@@
$\mathcal{F}\{\hat{\bm{x}}\odot\hat{\bm{a}}$=($\}=(\mathcal{F}\{\hat{\bm{x}}\}*\mathcal{F}\{\hat{\bm{a}}\})[k]$ (10)
@@
Figure 2: Log amplitude spectra of $\hat{\bm{x}}$ (blue), $\hat{\bm{a}}$ (green), and $\hat{\bm{x}}\odot\hat{\bm{a}}$ (red) are shown for a 60 Hz sinusoidal signal $\bm{x}$ when the ReLU [25] function is applied. The sampling frequency of the original signal $\bm{x}$ is 1 kHz. The dotted lines represent the frequencies of the harmonic components present in Eq. (1). Aliasing is still evident, even when applying the anti-aliasing nonlinear operation [19].
@@ III-C1 Frequency-domain convolution
($(\mathcal{F}\{\bm{x}\}*\bm{l})[k]=\sum_{m=0}^{L-1}\mathcal{F}\{\bm{x}\}[\,(k+m)~{}\mathrm{mod}~{}N]~{}\bm{l}[m].$ (12)
@@ IV Proposed Method
Figure 3: This diagram provides an overview of Wavehax. The kernel width of the 1D convolution is set to 7, while the kernel size of the depthwise convolution is set to 7 $\times$ 7. The numbers of hidden channels, denoted as $C$ and $C^{\prime}$, are set to 32 and 64, respectively. The number of frequency bins, $F$, is set to 241, calculated as half of the discrete Fourier transform points plus one. $T$ and $N$ represent the number of time steps in the waveforms and time frames in the features, respectively.
@@ IV-A Harmonic prior
Analogous to prior distributions in Bayesian statistics, we use the term $prior$ to refer to signals considered by the model before observing any training examples. Intuitively, the prior signal acts as an explicit bias, guiding the model in generating the output waveform [^2] Note that GAN-based vocoders, such as PWG [14], can be seen as using a non-informative prior, where Gaussian noise serves as input. This absence of prior knowledge forces the models to infer the spectral and temporal structures of output waveforms entirely from the training data.. Periodic priors have proven effective for explicit prosodic control in numerous studies, particularly in GAN vocoders [30, 53, 54, 55, 56, 57, 58], as well as in other frameworks [47, 59, 60]. We further analyze and provide a theoretical explanation of how periodic priors enhance neural vocoders in both the time and time-frequency domains.
@@ IV-A1 Periodic prior for time domain models
Our hypothesis is that time-domain neural vocoders can effectively generate harmonic components from periodic priors by modulating them through nonlinear operations. To substantiate this hypothesis, we first examine Taylor expansion of a general pointwise nonlinear function, $f:\mathbb{R}\rightarrow\mathbb{R}$, assuming this function satisfies the condition for Taylor expansion. Applying the Taylor series of $f$ to a time-domain signal $\bm{x}$ defined at each time step $n$, we obtain the following expression for the transformed signal $f(\bm{x})[n]=f(\bm{x}[n])$:
@@
This expansion shows that $f(\bm{x})$ is a weighted sum of powers of $\bm{x}^{k}$, and the magnitude of the derivatives $f^{(k)}$ influence the degree of aliasing. According to the inverse convolution theorem, the powers of a signal in the time domain correspond to repeated convolutions in the frequency domain, implying that higher-order terms introduce additional frequency components. Figure 4 illustrates the frequency characteristics of each power up to the 6th order when $\bm{x}$ is a sinusoidal signal.
@@ Appendix A Harmonic Decomposition of Sine Powers
$\mathrm{sin}$ (${}^{k+1}(\theta)=\mathrm{sin}(\theta)\sum_{m=0}^{k}\bigl{\{}a_{m}\mathrm{sin}(m\theta)+b_{m}\mathrm{cos}(m\theta)\bigl{\}}$

### v2

@@ II-D Time-frequency domain neural vocoders
HiNet [46] hierarchically estimates the log amplitude and phase spectrogram, ultimately reconstructing the waveform using inverse STFT (iSTFT). It estimates a log amplitude spectrogram using a feed-forward network and conditions a neural source-filter (NSF) [47] model with the log amplitude spectrogram to generate a time-domain waveform, from which the phase spectrogram is extracted via STFT. The excitation generation module of the NSF constructs harmonic signals by accumulating phase rotation angles for each time step, determined by $\text{F}_{0}$. This process ensures a coherent phase structure of the output waveforms, aiding efficient phase spectrogram estimation and allowing HiNet to sidestep the challenges of direct phase estimation. However, the phase spectrogram derived from time-domain processing involving nonlinear operations can introduce aliasing.
@@ III-A1 Time-domain convolution
[…]
@@ III-B2 Equivalent interpretation and implementation
[…]
@@
$\mathcal{F}\{\hat{\bm{x}}\odot\hat{\bm{a}}$ $\}=(\mathcal{F}\{\hat{\bm{x}}\}*\mathcal{F}\{\hat{\bm{a}}\})[k]$ (10)
@@
Figure 2: Amplitude spectra in dB of $\hat{\bm{x}}$ (blue), $\hat{\bm{a}}$ (green), and $\hat{\bm{x}}\odot\hat{\bm{a}}$ (red) are shown for a 60 Hz sinusoidal signal $\bm{x}$ when the ReLU [25] function is applied. The sampling frequency of the original signal $\bm{x}$ is 1 kHz. The dotted lines mark expected harmonic frequencies from Eq. (1). Despite the use of an anti-aliasing nonlinear operation [19], aliasing artifacts are still clearly observed in the green and red spectra.
@@ III-C1 Frequency-domain convolution
$(\mathcal{F}\{\bm{x}\}*\bm{l})[k]=\sum_{m=0}^{L-1}\mathcal{F}\{\bm{x}\}[\,(k+m)~\mathrm{mod}~N]~\bm{l}[m].$ (12)
@@ IV Proposed Method
Figure 3: This diagram provides an overview of Wavehax. The kernel width of the 1D convolution is set to 7, while the kernel size of the depthwise convolution is set to 7 $\times$ 7. The numbers of hidden channels, denoted as $C$ and $C^{\prime}$, are set to 32 and 64, respectively. The number of frequency bins, $F$, is set to 241, calculated as half of the discrete Fourier transform points plus one. $T$ and NN represent the number of time steps in the waveforms and time frames in the features, respectively.
@@ IV-A Harmonic prior
Analogous to prior distributions in Bayesian statistics, we use the term priorprior to refer to signals considered by the model before observing any training examples. Intuitively, the prior signal acts as an explicit bias, guiding the model in generating the output waveform [^2] Note that GAN-based vocoders, such as PWG [14], can be seen as using a non-informative prior, where Gaussian noise serves as input. This absence of prior knowledge forces the models to infer the spectral and temporal structures of output waveforms entirely from the training data.. Periodic priors have proven effective for explicit prosodic control in numerous studies, particularly in GAN vocoders [30, 53, 54, 55, 56, 57, 58], as well as in other frameworks [47, 59, 60]. We further analyze and provide a theoretical explanation of how periodic priors enhance neural vocoders in both the time and time-frequency domains.
@@ IV-A1 Periodic prior for time domain models
Our hypothesis is that time-domain neural vocoders can effectively generate harmonic components from periodic priors by modulating them through nonlinear operations. To substantiate this hypothesis, we first examine Maclaurin expansion of a general pointwise nonlinear function, $f:\mathbb{R}\rightarrow\mathbb{R}$, assuming this function satisfies the condition for Maclaurin expansion. Applying the Maclaurin series of $f$ to a time-domain signal $\bm{x}$ defined at each time step $n$, we obtain the following expression for the transformed signal $f(\bm{x})[n]=f(\bm{x}[n])$:
@@
This expansion shows that $f(\bm{x})$ is a weighted sum of powers of $\bm{x}^{k}$, and the magnitude of the derivatives $f^{(k)}$ influences the degree of aliasing. According to the inverse convolution theorem, the powers of a signal in the time domain correspond to repeated convolutions in the frequency domain, implying that higher-order terms introduce additional frequency components. Figure 4 illustrates the frequency characteristics of each power up to the 6th order when $\bm{x}$ is a sinusoidal signal.
@@ Appendix A Harmonic Decomposition of Sine Powers
$\mathrm{sin}$ $^{k+1}(\theta)=\mathrm{sin}(\theta)\sum_{m=0}^{k}\bigl\{a_{m}\mathrm{sin}(m\theta)+b_{m}\mathrm{cos}(m\theta)\bigl\}$